OpenStudy (anonymous):

Which expression is a cube root of -1+i√3? A. 3√2(cos40degrees+isin40degrees) B. 3√2(cos120degrees+isin120degrees) C. 3√4(cos280degrees+isin280degrees) D. 3√4(cos320degrees+isin320degrees)

OpenStudy (kropot72):

Put the expression into polar form:\[-1+i \sqrt{3}=2(\cos 120+i \sin 120)\] By Demoivre's Theorem: \[\sqrt[3]{2(\cos 120+i \sin 120}=\sqrt[3]{2}(\cos 40+i \sin 40)\] The correct choice is A.

OpenStudy (anonymous):

Okay, great! Thank you!

OpenStudy (kropot72):

You're welcome :)